Each element of a circular conical pile of sand 6ft. high is inclined 45 degrees to the horizontal. How many cubic feet of sand does the pile contain?


1
Expert's answer
2021-12-12T17:51:52-0500

Volume of a circular cone is;

"πr^2\frac{h}{3}"


"r=6tan45\u00b0=6"


"\therefore V=π×6^2\times\frac{6}{3}"


"=226.29" cubic feet